Two cards are drawn successively with replacement from a pack of $52$ cards. The probability of drawing two aces is

  • A
    $\frac{1}{169}$
  • B
    $\frac{1}{221}$
  • C
    $\frac{1}{2652}$
  • D
    $\frac{4}{663}$

Explore More

Similar Questions

$A$ and $B$ are independent events. Their probabilities are $3/10$ and $2/5$ respectively. What is the probability that exactly one of the events occurs?

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

If $A$ and $B$ are two independent events and $P(A)=\frac{3}{5}$ and $P(B)=\frac{2}{3}$,then $P(A' \cap B')=$

$A$ bag $B$ contains $4$ white balls and $2$ black balls. Another bag $C$ contains $3$ white balls and $5$ black balls. If one ball is drawn randomly from each bag, then the probability that the two balls drawn are both white is

$A$ purse contains $4$ copper coins and $3$ silver coins. $A$ second purse contains $6$ copper coins and $4$ silver coins. $A$ purse is chosen randomly and a coin is taken out of it. What is the probability that it is a copper coin?

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo